12a.A right triangle has a hypotenuse of length 2.00 m, and one of its angles is 35.0°. What are the lengths of the following sides?

(a) the side opposite the 35.0° angle

 


(b) the side adjacent to the 35.0° angle

 


1
Expert's answer
2022-06-06T08:54:29-0400

(a)

"a=c\sin\theta=2.00*\sin 35.0^\circ=1.15\:\rm m"

(b)

"b=c\cos\theta=2.00*\cos 35.0^\circ=1.64\:\rm m"